# Deep Hedging with Options

### v2

@@ A.1   Details for the MSGD training approach
The MSGD method estimates the objective function $\mathcal{O}_{\lambda}(\theta,\,l)$ by using small samples of the hedging error, referred to as batches. Let $\mathbb{B}_{j}=\left\{\xi_{T,i}^{(\tilde{\phi}_{\theta_{j}},l_{j})}\right\}_{i=1}^{B_{\mbox{\scriptsize{batch}}}}$ be the $j$-th batch simulated with policy parameters $\theta_{j}$ and $l_{j}$. Using a subset from generated paths, it represents a set of hedging errors
$\xi_{T,i}^{(\tilde{\phi}_{\theta_{j}},l_{j})}=\Psi(S^{(j)}_{T,i})-V_{T,i}^{(\tilde{\phi}_{\theta_{j}},l_{j})}\quad\mbox{for}\quad i\in\{1,\dots,B_{\mbox{\scriptsize{batch}}}\},\,j\in\{1,\dots,N_{\mbox{\scriptsize{batch}}}\},$
@@
where $S^{(j)}_{T,i}$ and $V_{T,i}^{(\tilde{\phi}_{\theta_{j}},l_{j})}$ respectively represent the time-$T$ underlying asset price and the terminal value of the hedging portfolio for path $i$ of batch $j$. The batch size is $B_{\mbox{\scriptsize{batch}}}=1000$, and the total number of batches is $N_{\mbox{\scriptsize{batch}}}=400$. The objective function estimates for batch $\mathbb{B}_{j}$ are
@@
$\hat{\mathcal{O}}_{\lambda}^{\left(\mbox{\scriptsize{MSE}}\right)}(\theta_{j},l_{j},\mathbb{B}_{j})$ $=\frac{1}{B_{\mbox{\scriptsize{batch}}}}\sum_{i=1}^{B_{\mbox{\scriptsize{batch}}}}\left(\xi_{T,i}^{(\tilde{\phi}_{\theta_{j}},l_{j})}\right)^{2}+\lambda\cdot\widehat{SC}(\theta_{j},l_{j},\mathbb{B}_{j}),$
@@
$\hat{\mathcal{O}}_{\lambda}^{\left(\mbox{\scriptsize{SMSE}}\right)}(\theta_{j},l_{j},\mathbb{B}_{j})$ $=\frac{1}{B_{\mbox{\scriptsize{batch}}}}\sum_{i=1}^{B_{\mbox{\scriptsize{batch}}}}\left(\xi_{T,i}^{(\tilde{\phi}_{\theta_{j}},l_{j})}\right)^{2}\mathbbm{1}_{\left\{\xi_{T,i}^{(\tilde{\phi}_{\theta_{j}},l_{j})}\geq 0\right\}}+\lambda\cdot\widehat{SC}(\theta_{j},l_{j},\mathbb{B}_{j}),$
@@
$\hat{\mathcal{O}}_{\lambda}^{\left(\mbox{\scriptsize{CVaR}}\right)}(\theta_{j},l_{j},\mathbb{B}_{j})$ $=\widehat{\mbox{VaR}}_{\alpha}(\mathbb{B}_{j})+\frac{1}{(1-\alpha)B_{\mbox{\scriptsize{batch}}}}\sum_{i=1}^{B_{\mbox{\scriptsize{batch}}}}\max\left(\xi_{T,i}^{(\tilde{\phi}_{\theta_{j}},l_{j})}-\widehat{\mbox{VaR}}_{\alpha}(\mathbb{B}_{j}),0\right)$
$\quad+\lambda\cdot\widehat{SC}(\theta_{j},l_{j},\mathbb{B}_{j}),$
@@
$\widehat{SC}(\theta_{j},l_{j},\mathbb{B}_{j})=\frac{1}{B_{\mbox{\scriptsize{batch}}}}\sum_{i=1}^{B_{\mbox{\scriptsize{batch}}}}\mathbbm{1}_{\left\{\max_{t\in\{0,\dots,T\}}\left[P_{t,i}-V_{t,i}^{(\tilde{\phi}_{\theta_{j}},l_{j})}\right]>V_{0,i}^{(\tilde{\phi}_{\theta_{j}},l_{j})}\right\}},$
[…]
Appendix B Joint implied volatility and return model
B.1   Daily implied volatility surface
@@
B.2   Joint implied volatility and return dynamics
@@ C.1   Leland model
C.2   Delta-gamma hedging
@@
where $\Delta^{\mathcal{P}}_{t}$, $\Gamma^{\mathcal{P}}_{t}$, and $\Delta^{(\text{O})}_{t}$, $\Gamma_{t}^{(\text{O})}$ represent the Black-Scholes delta and gamma of the hedged portfolio and the hedging option, respectively. For all Black-Scholes Greeks we use the implied volatility $\sigma(M_{t},\tau_{t},\beta_{t})$ from the static surface as the volatility input parameter.
Appendix D Soft constraint regularization
Figure 11: Risk measure and soft constraint values.
  • Results are computed over 100,000 out-of-sample paths according to the conditions outlined in Section 4.3.1. The hedge consists of an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days.
@@ Appendix E Impact of state variable inclusion on hedging performance
$\mathcal{S}\backslash\{\mathcal{P}_{t},\Delta_{t}^{P},\gamma_{t}^{P}\}$ 0.195 0.089 0.696
$\mathcal{S}\backslash\{\mathcal{P}_{t}\}$ 0.128 0.069 0.680
@@ Appendix F Statistical arbitrage
[…]
@@
  • Results are computed over 100,000 out-of-sample paths according to the conditions outlined in Section 4.3.1. The hedge consists of an ATM straddle with a maturity of $T=63$ days. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days. The transaction cost for the underlying asset is set to $\kappa_{1}=0.05\%$, except for the first column where $\kappa_{1}=0\%$.
@@
Online Appendix (not part of the paper)
@@ Appendix G Systematic outperformance of RL agents
  • Results are computed using bootstrapping with a sample size of 1,000 over 100,000 out-of-sample paths according to the conditions outlined in Section 4.3.1. The hedge consists of an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days. Transaction cost levels are set to 0.05% for $\kappa_{1}$ and 0.5% for $\kappa_{2}$.

### v3

@@ Appendix B Details for the MSGD training approach
The MSGD method estimates the objective function $\mathcal{O}(\theta;\,\lambda)$ by using small samples of the hedging error, referred to as batches. Let $\mathbb{B}_{j}=\left\{\xi_{T,i}^{\tilde{\phi}_{\theta_{j}}}\right\}_{i=1}^{B_{\mbox{\scriptsize{batch}}}}$ be the $j$-th batch simulated with policy parameters $\theta_{j}$. Using a subset from generated paths, it represents a set of hedging errors
$\xi_{T,i}^{\tilde{\phi}_{\theta_{j}}}=\Psi(S^{(j)}_{T,i})-V_{T,i}^{\tilde{\phi}_{\theta_{j}}}\quad\mbox{for}\quad i\in\{1,\dots,B_{\mbox{\scriptsize{batch}}}\},\,j\in\{1,\dots,N_{\mbox{\scriptsize{batch}}}\},$
@@
where $S^{(j)}_{T,i}$ and $V_{T,i}^{\tilde{\phi}_{\theta_{j}}}$ respectively represent the time-$T$ underlying asset price and the terminal value of the hedging portfolio for path $i$ of batch $j$. The batch size is $B_{\mbox{\scriptsize{batch}}}=1000$, and the total number of batches is $N_{\mbox{\scriptsize{batch}}}=400$. The objective function estimates for batch $\mathbb{B}_{j}$ are
@@
$\hat{\mathcal{O}}^{\left(\mbox{\scriptsize{MSE}}\right)}(\theta_{j};\lambda,\mathbb{B}_{j})$ $=\frac{1}{B_{\mbox{\scriptsize{batch}}}}\sum_{i=1}^{B_{\mbox{\scriptsize{batch}}}}\left(\xi_{T,i}^{\tilde{\phi}_{\theta_{j}}}\right)^{2}+\lambda\cdot\widehat{SC}(\theta_{j},\mathbb{B}_{j}),$
@@
$\hat{\mathcal{O}}^{\left(\mbox{\scriptsize{SMSE}}\right)}(\theta_{j};\lambda,\mathbb{B}_{j})$ $=\frac{1}{B_{\mbox{\scriptsize{batch}}}}\sum_{i=1}^{B_{\mbox{\scriptsize{batch}}}}\left(\xi_{T,i}^{\tilde{\phi}_{\theta_{j}}}\right)^{2}\mathbbm{1}_{\left\{\xi_{T,i}^{\tilde{\phi}_{\theta_{j}}}\geq 0\right\}}+\lambda\cdot\widehat{SC}(\theta_{j},\mathbb{B}_{j}),$
@@
$\hat{\mathcal{O}}^{\left(\mbox{\scriptsize{CVaR}}\right)}(\theta_{j};\lambda,\mathbb{B}_{j})$ $=\widehat{\mbox{VaR}}_{\alpha}(\mathbb{B}_{j})+\frac{1}{(1-\alpha)B_{\mbox{\scriptsize{batch}}}}\sum_{i=1}^{B_{\mbox{\scriptsize{batch}}}}\max\left(\xi_{T,i}^{\tilde{\phi}_{\theta_{j}}}-\widehat{\mbox{VaR}}_{\alpha}(\mathbb{B}_{j}),0\right)+\lambda\cdot\widehat{SC}(\theta_{j},\mathbb{B}_{j}),$
@@
$\widehat{SC}(\theta_{j},\mathbb{B}_{j})=\frac{1}{B_{\mbox{\scriptsize{batch}}}}\sum_{i=1}^{B_{\mbox{\scriptsize{batch}}}}\mathbbm{1}_{\left\{\max_{t\in\{0,\dots,T\}}\left[P_{t,i}-V_{t,i}^{\tilde{\phi}_{\theta_{j}}}\right]>V_{0,i}^{\tilde{\phi}_{\theta_{j}}}\right\}},$
and $\widehat{\mbox{VaR}}_{\alpha}(\mathbb{B}_{j})=\xi_{T,\left[\lceil\alpha\cdot B_{\mbox{\scriptsize{batch}}}\rceil\right]}^{\tilde{\phi}_{\theta_{j}}}$ is the value-at-risk estimation derived from the ordered sample $\left\{\xi_{T,[i]}^{\tilde{\phi}_{\theta_{j}}}\right\}_{i=1}^{B_{\mbox{\scriptsize{batch}}}}$, where $\lceil\cdot\rceil$ is the ceiling function. These empirical approximations are used to estimate the gradient of the objective function required in Equation (7). The gradient of these empirical objective functions has analytical expressions for FFNN, LSTM and RNN-FNN networks, which can be computed through backpropagation, see for instance Goodfellow et al., (2016).
Appendix C Joint implied volatility and return model
C.1  Daily implied volatility surface
@@
C.2  Joint implied volatility and return dynamics
@@ D.1  Leland model
D.2  Delta-gamma hedging
@@
where $\Delta^{\mathcal{P}}_{t}$, $\Gamma^{\mathcal{P}}_{t}$, and $\Delta^{(\text{O})}_{t}$, $\Gamma_{t}^{(\text{O})}$ represent the delta and gamma of the hedged portfolio and of the hedging option, respectively. The self-financing constraint (1) fully determines $\phi^{(r)}_{t+1}$. For all Greeks we use the implied volatility $\sigma(M_{t},\tau_{t},\beta_{t})$ from the static surface as the volatility input parameter.
D.3  No-trade region
This is a recursive construction. In the time interval (t1,t](t-1,t], denote the hedging portfolio with the no-trade threshold $\ell$ by $\phi_{t}^{(\ell)}=\left(\phi^{(\ell,r)}_{t},\phi^{(\ell,S)}_{t},\phi^{(\ell,\text{O})}_{t}\right)$. At time $t$, its value is
$V^{(\ell,\phi)}_{t}=\phi^{(\ell,r)}_{t}\text{e}^{r_{t}\Delta}+\phi^{(\ell,S)}_{t}S_{t}\text{e}^{q_{t}\Delta}+\phi^{(\ell,\text{O})}_{t}O_{t}\left(T^{*}\right).$
The no-trade region constraint is set up such that
$\left(\phi^{(\ell,S)}_{t+1},\phi^{(\ell,\text{O})}_{t+1}\right)$ $=\begin{cases}\left(\phi^{(\ell,S)}_{t},\phi^{(\ell,\text{O})}_{t}\right),&\text{if }\left|\phi_{t}^{(\ell,S)}-\left(\Delta^{\mathcal{P}}_{t}-\frac{\Gamma^{\mathcal{P}}_{t}}{\Gamma_{t}^{(\text{O})}}\Delta^{(\text{O})}_{t}\right)\right|+\left|\phi^{(\ell,\text{O})}_{t}-\frac{\Gamma^{\mathcal{P}}_{t}}{\Gamma_{t}^{(\text{O})}}\right|\leq\ell,\\[4.0pt] \left(\Delta^{\mathcal{P}}_{t}-\frac{\Gamma^{\mathcal{P}}_{t}}{\Gamma_{t}^{(\text{O})}}\Delta^{(\text{O})}_{t},\,\frac{\Gamma^{\mathcal{P}}_{t}}{\Gamma_{t}^{(\text{O})}}\right),&\text{otherwise}.\end{cases}$
The bank account position is
$\phi^{(\ell,r)}_{t+1}=V^{(\ell,\phi)}_{t}-\phi^{(\ell,S)}_{t+1}S_{t}-\phi^{(\ell,O)}_{t+1}O_{t}(T^{*})-\kappa_{1}\left|\phi^{(\ell,S)}_{t+1}-\phi^{(\ell,S)}_{t}\right|S_{t}-\kappa_{2}\left|\phi^{(\ell,O)}_{t+1}-\phi^{(\ell,O)}_{t}\right|O_{t}(T^{*}).$
The parameter $\ell$ is optimized by minimizing one of the three objective functions computed on the entire learning set:
$\hat{\mathcal{O}}^{\left(\mbox{\scriptsize{MSE}}\right)}(\ell)$ $=\frac{1}{N}\sum_{i=1}^{N}\left(\xi_{T,i}^{\phi^{(\ell)}}\right)^{2}$
$\hat{\mathcal{O}}^{\left(\mbox{\scriptsize{SMSE}}\right)}(\ell)$ $=\frac{1}{N}\sum_{i=1}^{N}\left(\xi_{T,i}^{\phi^{(\ell)}}\right)^{2}\mathbbm{1}_{\left\{\xi_{T,i}^{\phi^{(\ell)}}\geq 0\right\}}$
$\hat{\mathcal{O}}^{\left(\mbox{\scriptsize{CVaR}}\right)}(\ell)$ $=\widehat{\mbox{VaR}}_{\alpha}+\frac{1}{(1-\alpha)N}\sum_{i=1}^{N}\max\left(\xi_{T,i}^{\phi^{(\ell)}}-\widehat{\mbox{VaR}}_{\alpha},0\right)$
where $\xi_{T,i}^{\phi^{(\ell)}}=\mathcal{P}_{T,i}-V_{T,i}^{\phi^{(\ell)}}$ and $\widehat{\mbox{VaR}}_{\alpha}=\xi_{T,\left[\lceil\alpha\cdot N\rceil\right]}^{\phi^{(\ell)}}$ is the value-at-risk estimate derived from the ordered sample $\left\{\xi_{T,[i]}^{\phi^{(\ell)}}\right\}_{i=1}^{N}$.
@@ Appendix E Impact of state variable inclusion on hedging performance
[…]
@@ Appendix F Statistical arbitrage
Risk measure $\kappa_{1}=\kappa_{2}=0\%$ $\kappa_{2}$=0.5% $\kappa_{2}$=1% $\kappa_{2}$=1.5% $\kappa_{2}$=2%
@@
  • Results are computed over 100,000 out-of-sample paths according to the conditions outlined in Section˜4.3.1. The hedge consists of an ATM straddle with a maturity of $T=63$ days. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days. The transaction cost for the underlying asset is set to $\kappa_{1}=0.05\%$, except for the first column where $\kappa_{1}=0\%$.
@@
Supplementary material (not part of the paper)
@@ Appendix G Systematic outperformance of RL agents
  • Results are computed using bootstrapping with a sample size of 1,000 over 100,000 out-of-sample paths according to the conditions outlined in Section˜4.3.1. The hedge consists of an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days. Transaction cost levels are set to 0.05% for $\kappa_{1}$ and 0.5% for $\kappa_{2}$.
@@ Appendix H JIVR Model parameters
Appendix I Impact of no-trade regions
The hedging cost
$\mbox{HC}_{l}=\sum_{t=0}^{T-1}e^{-r\Delta t}\mathcal{HC}_{t},$ (13)
  • Results are computed over 100,000 out-of-sample paths according to the conditions outlined in Section˜4.3.
Conversely, as $\kappa_{2}$ increases, RL agents retain high rebalancing frequency, but keep average transaction costs to a level similar to DG. Thus, more gradual and frequent adjustments from RL mitigate risk more effectively than DG as documented in Section˜4.4.2, while leading to similar transaction costs.
Appendix J Soft constraint regularization
Figure 17: Risk measure and soft constraint values.
  • Results are computed over 100,000 out-of-sample paths according to the conditions outlined in Section˜4.3.1. The hedge consists of an ATM straddle with a maturity of $T=63$ days and an average value of \$7.55. The hedging instrument is an ATM call option with a maturity of $T^{*}=84$ days.
Appendix K In-sample backtest
  • Results are computed based on the observed P&L from hedging 296 straddle positions with maturity 63-days under real market conditions observed from May 1, 1996, to December 31, 2020. A new ATM straddle is considered every 21 business days. Agents are trained according to the conditions outlined in Section˜4.3 using an ATM call option with a maturity of $T^{*}=84$ days as the hedging instrument.